# Explicit Constructions of Sum-Rank Metric Codes from Quadratic Kummer Extensions

### v1

@@ III A General Setting of Sum-Rank Metric Codes
In this section, we present a general construction of sum-rank metric codes through function field extensions, which differs from the approach by E. Berardini and X. Caruso in [4].
@@ III-A The Main Idea of Matrices over $\mathbb{F}_{q}$
Let $F/\mathbb{F}_{q}$ be a function field and $F^{\prime}/\mathbb{F}_{q^{m}}$ a Galois extension of $F/\mathbb{F}_{q}$ with $[F^{\prime}:F]=n$. Suppose that $P$ is a place of $F$ and $t\in F$ is a prime element of $P$. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ is given by
@@
The quotient $V:=\mathcal{O}^{\prime}_{P}/t\mathcal{O}^{\prime}_{P}$ forms a vector space over $\mathcal{O}_{P}/P$ with basis
@@
For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and place $P^{\prime}$ with $P^{\prime}|P$, observe that for $z\in\mathcal{O}^{\prime}_{P}$:
@@
since $\sigma^{-1}(P^{\prime})|P$. This implies $\sigma(z)\in\mathcal{O}^{\prime}_{P}$, hence $\sigma(V)=V$. The set
@@
forms another basis of $V$ over $\mathcal{O}_{P}/P$. Then $\sigma$ induces a matrix $\bar{A}_{\sigma}$ over $\mathcal{O}_{P}/P$ with respect to $\{z_{i}\}$.
Similarly, for $f\in\mathcal{O}^{\prime}_{P}$, consider the multiplication $\mu_{f}:\mathcal{O}^{\prime}_{P}\to\mathcal{O}^{\prime}_{P}$ defined by $\mu_{f}(z)=fz$. This induces an $\mathcal{O}_{P}/P$-linear map $\bar{\mu}_{f}:V\to V$ via
@@
This induces a matrix $\bar{B}_{f}$ with entries in $\mathcal{O}_{P}/P$.
@@
The matrix $A=(a_{ij})$ associated to $\sigma$ has entries in $\mathcal{O}_{P}$ due to the integral basis property. Under the reduction map
$\pi:\mathcal{O}_{P}\to\mathcal{O}_{P}/P,$
we obtain $\bar{A}=(\pi(a_{ij}))$. This explains the notation $\bar{A}_{\sigma}$ and $\bar{B}_{f}$.
When $P$ is rational (i.e., $\mathcal{O}_{P}/P=\mathbb{F}_{q}$), both $\sigma\in\text{Gal}l(F^{\prime}/F)$ and $f\in\mathcal{O}^{\prime}_{P}$ induce matrices over $\mathbb{F}_{q}$. Let $\text{Gal}(F^{\prime}/F)=\{\sigma_{1},\ldots,\sigma_{n}\}$, and consider functions $f_{1},\ldots,f_{n}\in\mathcal{O}^{\prime}_{P}$. We define a linear operator
@@
This $F^{\prime}$-linear operator naturally induces a linear morphism on $F$. We further define a map $\varepsilon_{P}$ satisfying
@@
where $B_{i}$ and $A_{i}$ are matrices representing $\mu_{f_{i}}$ and $\sigma_{i}$ for $1\leq i\leq n$, respectively. Furthermore, let
$\bar{\varepsilon}_{P}(L)=\sum\limits_{i=1}^{n}\bar{A_{i}}\bar{B_{i}}=(\pi(M_{ij})).$
We establish the following lemma.
@@
Thus $f\sigma(g)=[g_{1},\ldots,g_{n}]B_{f}A_{\sigma}[z_{1},\ldots,z_{n}]^{T}$. The conclusion follows from the linearity of $\mu_{f}$ and $\sigma$.
@@
Any solution $[g_{1},\ldots,g_{n}]$ satisfies $L(g_{1}z_{1}+\cdots+g_{n}z_{n})=0$. Therefore ${\rm ker}(L)$ contains non-zero elements if and only if the system has non-trivial solutions, which occurs precisely when $\det(\varepsilon_{P}(L))=0$. ∎
III-B Sum-rank Codes in General Case
Now we present the construction of sum-rank metric codes. Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$. Assume that the divisor $Q_{0}$ is totally ramified in $F^{\prime}/F$ with $P_{0}|Q_{0}$. Let $G=nQ_{0}$, $G^{\prime}=nP_{0}$ and $D=Q_{1}+\cdots+Q_{s}$. Define the evaluation map:
$\text{ev}_{D,1}:\bigoplus_{i=1}^{n}\mathcal{L}(G^{\prime})$ $\to\Pi$
$L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$$\mapsto\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))\}$
Since $f_{1},\ldots,f_{n}$ have poles only at $P_{0}$, they belong to $\mathcal{O}^{\prime}_{Q_{i}}$ for all $1\leq i\leq s$. Thus $\varepsilon_{Q_{i}}(L)$ are well-defined. Moreover, the function space $\mathcal{L}(G^{\prime})$ forms an $\mathbb{F}_{q}$-linear space and $L$ is $F$-linear. It follows that $\text{ev}_{D,1}$ is an $\mathbb{F}_{q}$-linear map. Thus the image of $\text{ev}_{D,1}$ is a subspace of $\Pi$. The code $C_{sr,1}(D,G)$ is defined as the image of $\text{ev}_{D,1}$.
@@ III-B Sum-rank Codes in General Case
The code $C_{sr,1}$ resembles the code $C(1,G,Q_{1},\ldots,Q_{s})$ defined in [4, Definition 4]. In their work, E. Berardini and X. Caruso constructed codes through the isomorphism chain
@@
where $F_{i}$ denotes the completion of $F$ at $Q_{i}$, while $F^{\prime}_{i}$ represents the product of completions of $F^{\prime}$ at places lying over $Q_{i}$ for each $i=1,\ldots,n$. Our approach directly employs $L$ to express linearized-polynomial-like maps, corresponding to elements of $D_{F^{\prime},1}$. We subsequently utilize the integral closure $\mathcal{O}^{\prime}_{Q_{i}}$ to construct explicit matrices $\bar{\varepsilon}_{Q_{i}}(L)$, where $\varepsilon_{Q_{i}}(L)$ is induced by $L$ at $Q_{i}$.
Furthermore, while previous work proposed several hypotheses for code parameter determination, the ring $D_{F^{\prime}_{i},1}$ fails to satisfy all necessary conditions. In subsequent sections, we provide more explicit dimension and minimum distance bounds by considering subspaces of $\bigoplus_{i=1}^{n}\mathcal{L}(G^{\prime})$.
@@
In the construction of $C_{sr,1}$, each rational place $Q\neq Q_{0}$ yields a well-defined $n\times n$ matrix $\bar{\varepsilon}_{P}(L)$ over $\mathbb{F}_{q}$. By the Hasse-Weil bound, the function field $F/\mathbb{F}_{q}$ contains at most $q+1+2g(F)\sqrt{q}$ rational places, implying $s\leq q+2g(F)\sqrt{q}$. ∎
@@ IV Sum-Rank Metric Codes via Quadratic Extension of Algebraic Function Fields
IV-A Parameters of the General Construction
We first establish some necessary lemmas before deriving the dimension and minimum sum-rank distance for $C_{sr,1}$.
@@ IV-A Parameters of the General Construction
Suppose that $Q\neq P$ is a rational place of $F$ with $Q$ totally ramified in $F^{\prime}/F$. Let $Q^{\prime}$ be the place of $F^{\prime}$ lying over $Q$. For $f_{1},\ldots,f_{n}\in\mathcal{L}(kQ^{\prime})$, we have
@@
Furthermore, if there exists $f_{i}$ satisfying
@@
Since $f_{i}\in\mathcal{L}(kQ^{\prime})$, the coefficients of $B_{f_{i}}$ are contained in $\mathcal{L}(kQ^{\prime})$ for $i=1,\ldots,n$. The discrete valuation of $\det(\varepsilon_{P}(L))$ at $Q^{\prime}$ therefore exceeds the minimal valuation in its determinant expansion. Consequently, we have $\det(\varepsilon_{P}(L))\in\mathcal{L}(knQ^{\prime})$. Since $Q$ is totally ramified, the conorm of $Q$ is
@@
By Lemma III.1 (ii) and the conorm properties, we have
@@
Note that for any $Q^{\prime}$ of $F^{\prime}$, there exists only one $Q$ in $F$ with $Q^{\prime}|Q$, thus $\det(\varepsilon_{P}(L))$ only has poles at $Q$. Therefore $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ)$. For any non-zero $g\in\mathcal{O}^{\prime}_{P}$, the equality
@@
holds for $i=1,\ldots,n$. W.l.o.g, assume that $v_{Q^{\prime}}(f_{1})<v_{Q^{\prime}}(f_{j})$ when $2\leq j\leq n$. This yields
@@
By the strict triangle inequality, we have
@@
Thus $L(g)\neq 0$ for $g\neq 0$. The conclusion follows from Lemma III.1 (iii). ∎
@@
Let $P$ be a place of $F$ with a prime element $t$. If $\det(\varepsilon_{P}(L))\neq 0$ for $L$ associated tp $Q\neq P$ while $rank(\bar{\varepsilon}_{P}(L))<n$, then $\det(\varepsilon_{P}(L))\in P$.
@@
The condition $rank(\bar{\varepsilon}_{P}(L))<n$ means $\det(\bar{\varepsilon}_{P_{\alpha}}(L))=0$. Let $\pi_{P}$ denote the residue class map at $P$. For $\varepsilon_{P}(L)=(M_{ij})$, we have

### v2

@@ III A General Setting of Sum-Rank Metric Codes
In this section, we introduce a general construction of sum-rank metric codes using function field extensions, which differs from the approach proposed by E. Berardini and X. Caruso in [4].
@@ III-A The Main Idea of Matrices over $\mathbb{F}_{q}$
Let $F/\mathbb{F}_{q}$ be a function field, and $F^{\prime}/\mathbb{F}_{q^{m}}$ be a Galois extension of $F/\mathbb{F}_{q}$ with $F^{\prime}:F]=n$. Suppose that $P$ is a place of $F$, and $t\in F$ is a prime element of $P$. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ is given by
@@
The quotient $V:=\mathcal{O}^{\prime}_{P}/t\mathcal{O}^{\prime}_{P}$ forms a vector space over the finite field $\hat{F}_{P}$ with basis
@@
For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and a place $P^{\prime}$ with $P^{\prime}|P$, observe that for any $z\in\mathcal{O}^{\prime}_{P}$,
@@
since $\sigma^{-1}(P^{\prime})|P$. This implies $\sigma(z)\in\mathcal{O}^{\prime}_{P}$, and hence $\sigma(V)=V$. The set
@@
forms another basis of $V$ over $\hat{F}_{P}$. Then $\sigma$ induces an $n\times n$ matrix $\bar{A}_{\sigma}$ over $\hat{F}_{P}$ with respect to the basis $\{z_{i}\}$.
Similarly, for any $f\in\mathcal{O}^{\prime}_{P}$, consider the multiplication map $\mu_{f}:\mathcal{O}^{\prime}_{P}\to\mathcal{O}^{\prime}_{P}$ defined by $\mu_{f}(z)=fz$. This induces an $\hat{F}_{P}$-linear map $\bar{\mu}_{f}:V\to V$ via
@@
This also induces an $n\times n$ matrix $\bar{B}_{f}$ with entries in $\hat{F}_{P}$.
@@
The matrix $A=(a_{ij})$ associated with $\sigma$ has entries in $\mathcal{O}_{P}$ due to the integral basis property. Under the reduction map
$\pi:\mathcal{O}_{P}\to\hat{F}_{P},$
we obtain $\bar{A}=(\pi(a_{ij}))$. This justifies the notation $\bar{A}_{\sigma}$ and $\bar{B}_{f}$.
When $P$ is rational (i.e., $\hat{F}_{P}=\mathbb{F}_{q}$), both $\sigma\in\text{Gal}(F^{\prime}/F)$ and $f\in\mathcal{O}^{\prime}_{P}$ induce matrices over $\mathbb{F}_{q}$. Let $\text{Gal}(F^{\prime}/F)=\{\sigma_{1},\ldots,\sigma_{n}\}$, and consider functions $f_{1},\ldots,f_{n}\in\mathcal{O}^{\prime}_{P}$. We define a linear operator
@@
For any two functions $g_{1},g_{2}\in F$, we have
$L(g_{1}+g_{2})=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}(g_{1}+g_{2})=\sum\limits_{i=1}^{n}f_{i}(g_{1}+g_{2})=L(g_{1})+L(g_{2}).$
Thus, this $F^{\prime}$-linear operator naturally induces a linear morphism on $F$. We further define a map $\varepsilon_{P}$ such that
@@
where $B_{i}$ and $A_{i}$ are matrices representing $\mu_{f_{i}}$ and $\sigma_{i}$ for $1\leq i\leq n$, respectively. Then $\varepsilon_{P}(L)$ is an $n\times n$ matrix over $\mathcal{O}_{P}$. Moreover, let
$\bar{\varepsilon}_{P}(L)=\sum\limits_{i=1}^{n}\bar{A_{i}}\bar{B_{i}}=(\pi(M_{ij}))$
be the matrix over $\mathbb{F}_{q}$ induced by $\varepsilon_{P}(L)$. We now state the following lemma.
@@
Thus $f\sigma(g)=[g_{1},\ldots,g_{n}]B_{f}A_{\sigma}[z_{1},\ldots,z_{n}]^{T}$. The result follows from the linearity of the maps $\mu_{f}$ and $\sigma$.
@@
Any solution $g_{1},\ldots,g_{n}]$ satisfies $L(g_{1}z_{1}+\cdots+g_{n}z_{n})=0$. Therefore ${\rm ker}(L)$ contains nonzero elements if and only if the system has nontrivial solutions, which occurs precisely when $\det(\varepsilon_{P}(L))=0$. ∎
III-B Sum-rank Codes in the General Case
We now present the construction of sum-rank metric codes. Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$, and $Q_{0}$ be another place of $F$. Assume that $P_{0,1},\ldots,P_{0,m}$ are all places of $F^{\prime}$ lying over $Q_{0}$.Let $G=nQ_{0}$, $D=Q_{1}+\cdots+Q_{s}$, and $G_{j}$ are $n$ divisor of $F^{\prime}$ with
$\text{supp}(G_{j})\subset\{P_{0,1},\ldots,P_{0,m}\}.$
Define the evaluation map:
$\text{ev}_{D}:\mathcal{L}(G_{1})\oplus\cdots\oplus\mathcal{L}(G_{n})$ $\to\Pi=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{n\times n}$
$(f_{1},\ldots,f_{n})$ $\mapsto(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))$
where $L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$. Since $f_{1},\ldots,f_{n}$ have poles only at $P_{0}$, they belong to $\mathcal{O}^{\prime}_{Q_{i}}$ for all $1\leq i\leq s$. Thus $\varepsilon_{Q_{i}}(L)$ are well-defined. Note that $\mathcal{L}(G_{j})$ is an $\mathbb{F}_{q}$-linear space. Since $L$ is $F$-linear, it follows that $\text{ev}_{D}$ is an $\mathbb{F}_{q}$-linear map. Thus the image of $\text{ev}_{D}$ is a subspace of $\Pi$. The code $C_{sr}(D,G)$ is defined as the image of $\text{ev}_{D}$, which is exactly a linear space over $\mathbb{F}_{q}$.
@@ III-B Sum-rank Codes in the General Case
In the work of [4], E. Berardini and X. Caruso constructed codes through the isomorphism chain
@@
where $F_{i}$ denotes the completion of $F$ at $Q_{i}$, and $F^{\prime}_{i}$ denotes the product of completions of $F^{\prime}$ at places lying over $Q_{i}$ for each $i=1,\ldots,s$. The functor for generating the matrix in [4] has form $\sum\limits_{i=0}^{n-1}f_{i}\Phi^{i}$, where $\Phi^{r}=x$ for some $0\neq x\in F$, and $f_{i}\in\Lambda_{L,x}(G^{\prime})$. To determine the code parameters, E. Berardini and X. Caruso hypothesized that $D_{L,x}$ has no nonzero zero divisor.
Our approach directly employs the operator $L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$, corresponding to elements of $D_{F^{\prime},1}$ since $F^{\prime}/F$ is Galois. However, the ring $D_{F^{\prime},1}$ is not a domain because the polynomial $T^{r}-1\in F^{\prime}[T]$ is reducible. For the determination of code parameters, we consider $Q_{0}$ that is totally ramified or splits completely in $F^{\prime}/F$, and then present another choice for $f_{i}$, which is explicitly described in subsequent sections.
@@
In the construction of $C_{sr}$, each place $Q\neq Q_{0}$ yields a well-defined $n\times n$ matrix $\bar{\varepsilon}_{P}(L)$ over $\mathbb{F}_{q}$. By the Hasse-Weil bound, the function field $F/\mathbb{F}_{q}$ has at most $q+1+2g(F)\sqrt{q}$ rational places, which implies that $s\leq q+2g(F)\sqrt{q}+1$. ∎
@@ IV Sum-Rank Metric Codes via Quadratic Extension of Algebraic Function Fields
IV-A Construction with Totally Ramified Rational Places
We first establish some necessary lemmas before deriving the dimension and minimum sum-rank distance.
@@ IV-A Construction with Totally Ramified Rational Places
Let $P$ be a rational place of $F$. Suppose that $Q\neq P$ is a place of $F$ such that $Q$ is totally ramified in $F^{\prime}/F$. Let $Q^{\prime}$ be the place of $F^{\prime}$ lying over $Q$. For $f_{1},\ldots,f_{n}\in\mathcal{L}(kQ^{\prime})$, we have
@@
Furthermore, if there exists an $f_{i}$ satisfying
@@
Since $f_{i}\in\mathcal{L}(kQ^{\prime})$, the coefficients of $B_{f_{i}}$ belong to $\mathcal{L}(kQ^{\prime})$ for $i=1,\ldots,n$. Therefore, the discrete valuation of $\det(\varepsilon_{P}(L))$ at $Q^{\prime}$ is greater than the minimal valuation in its determinant expansion. Consequently, we have $\det(\varepsilon_{P}(L))\in\mathcal{L}(knQ^{\prime})$. As $Q$ is totally ramified, the conorm of $Q$ is
@@
By Lemma III.1 (ii) and the properties of the conorm, we obtain
@@
Note that for any place $Q^{\prime}$ of $F^{\prime}$, there is a unique place $Q$ in $F$ such that $Q^{\prime}|Q$, so $\det(\varepsilon_{P}(L))$ has poles only at $Q$. Hence, $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ)$. For any nonzero $g\in\mathcal{O}^{\prime}_{P}$, the equality
@@
holds for all $i=1,\ldots,n$. W.l.o.g, assume that $v_{Q^{\prime}}(f_{1})<v_{Q^{\prime}}(f_{j})$ for all $j=2,\ldots,n$. This implies that
@@
By the strict triangle inequality, it follows that
@@
Thus $L(g)\neq 0$ for any $g\neq 0$. The conclusion follows from Lemma III.1 (iii). ∎
@@
Let $P$ be a place of $F$ with a prime element $t$. If $\det(\varepsilon_{P}(L))\neq 0$ for $L$ associated to $Q\neq P$ while ${\rm rank}(\bar{\varepsilon}_{P}(L))<n$, then $\det(\varepsilon_{P}(L))\in P$.
@@
The condition ${\rm rank}(\bar{\varepsilon}_{P}(L))<n$ means that $\det(\bar{\varepsilon}_{P}(L))=0$. Let $\pi_{P}$ denote the residue class map at $P$. For $\varepsilon_{P}(L)=(M_{ij})$, we have
@@
Let $k,k_{1}$ be integers such that $k>k_{1}\geq 2g(F)-1$, and $P$ be a rational place of $F$. Then by the Riemann-Roch theorem, we obtain $\mathcal{L}(kP)$ and $\mathcal{L}(k_{1}P)$ are vector spaces over $\mathbb{F}_{q}$ with dimensions
$\ell_{k}:=k-g(F)+1,\ell_{k_{1}}:=k_{1}-g(F)+1,$
respectively. Let
$\mathcal{L}(kP)=\langle\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\rangle$
where $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k_{1}}}\}$ forms a basis for $\mathcal{L}(k_{1}P)$. We define
$\mathcal{L}^{*}_{k_{1}}(kP):=\langle\omega_{\ell_{k_{1}}+1},\ldots,\omega_{\ell_{k}}\rangle.$
It is obvious that $\mathcal{L}^{*}_{k_{1}}(kP)$ is a vector space over $\mathbb{F}_{q}$ of dimension $k-k_{1}$. Then we have:
Theorem IV.1.
Suppose that $F^{\prime}/F$ is a Galois extension of degree 2 with a totally ramified rational place $Q_{0}$ in $F^{\prime}/F$, and $Q^{\prime}_{0}|Q_{0}$. Let $s,k$ and $k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For distinct places $Q_{1},\ldots,Q_{s}$, the code
has parameters $4s,\ell_{k},d_{sr}\geq 2s-k]$.
Proof.
The length of $C_{sr,1}$ is exactly $4s$ over $\mathbb{F}_{q}$. Its dimension is determined by
Since $k<2s$, the evaluation map $\text{ev}_{D,1}$ is injective, implying that the dimension of $C_{sr,1}$ reduces to $k-k_{1}+\ell_{k_{1}}=\ell_{k}$.
Let $f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0})$ and $f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})$ with $f_{1}\neq 0$. Suppose that $rank(\bar{\varepsilon}_{Q_{i}}(L))=d_{i}$ for $i=1,\ldots,s$. Then the sum-rank weight of the codeword corresponding to $L$ is $d_{L}=\sum\limits_{i=1}^{s}d_{i}$. By Lemma IV.2, we have
$v_{Q_{i}}(\det(\varepsilon_{Q_{i}}(L)))\geq 2-d_{i}$
for each $i$. Furthermore, Lemma IV.1 shows that
$\det(\varepsilon_{Q_{i}}(L))\in\mathcal{L}(kQ_{0}).$
From $v_{Q^{\prime}_{0}}(f_{1})<v_{Q^{\prime}_{0}}(f_{2})$ and $f_{1}\neq 0$, it follows that $\det(\varepsilon_{Q_{i}}(L))\neq 0$.
Denote $F^{\prime}=F(y)$. Since $F^{\prime}:F]=2$, there exists a monic irreducible quadratic polynomial $\phi(T)\in F[T]$ that is the minimal polynomial of $y$ over $F$, write $\phi(T)=T^{2}+a_{1}T+a_{2}$.
When $f_{1}=0$, the matrix $\varepsilon_{Q_{i}}(L)$ takes the form
$A_{\sigma_{2}}\cdot\begin{bmatrix}&f_{21},&f_{22}\\ &-f_{22}a_{2},&f_{21}-f_{22}a_{1}\end{bmatrix}$
where $f_{2}=f_{21}z_{1}+f_{22}z_{2}$. Then $\det(\varepsilon_{Q_{i}}(L))=0$ means
$f_{21}^{2}-f_{21}f_{22}a_{1}+f_{22}^{2}a_{2}=0.$
If $f_{22}=0$, the previous equation implies $f_{21}=0$ and $f_{2}=0$. If $f_{21}\neq 0$, then the function $\frac{-f_{21}}{f_{22}}$ is a root of $\phi(T)$. Both cases lead to a contradiction. Thus $\det(\varepsilon_{Q_{i}}(L))$ cannot vanish, and it has at most $k$ zeros. This implies that
$\sum\limits_{i=1}^{s}(2-d_{i})\leq k.$
Consequently, the codeword satisfies $d_{L}\geq 2s-k$. Therefore, we have $d_{sr}\geq 2s-k$. ∎
IV-B Construction with Splitting Completely Rational Places
For a Galois extension $F^{\prime}/F$ of degree 2, the following lemmas are required to determine code parameters.